OpenStudy (anonymous):

The perimeter of a rectangle is 34 feet and the length is 3 feet longer than the width find the dimension , write the first equation 2x+2y=

OpenStudy (anonymous):

2x+2y = 34. let x be length and y be width. then x = y+3 Use x = y+3 in ^ equation and solve accordingly. :)

OpenStudy (anonymous):

So what's the length and the width in feet ? I don't get it !! Help please

OpenStudy (anonymous):

The thing is; 2x+2y = 34. 2(y+3) + 2y = 34. 2y+6+2y = 34 4y = 28 or y = 7. x = y+3 = 7+3 = 10. Width = 7 ft., length = 10 ft.

OpenStudy (anonymous):

Ok thanx I'm gonna review this so I understand it more